shown below is the financial date of company A and Company B at end of current year

company A company B

netsales(all on credit) $1440000 $ 1190000

cost of goods sold 1260000 825000

cash 36000 70000

Accounts receivable (Net) 180000 140000

inventory 504000 165000

current liabilities 240000 150000

Assume that the year end balance shown for accounts receivables and for inventory also represent the average balances of these accounts throughout the years

compute the following. (1) working capital (2) current ratio (3) Quick ratio (4) Number of times inventory turned over during the year and the average number of days required to turn over the inventory (5) Number of times accounts receivables turned over during the year and the average numbers of days required to collect accounts receivables (6) operating cycle


1
Expert's answer
2021-07-27T05:14:02-0400

(1) working capital = Current assets - Current liabilities

Company A

"(36000+180000+504000)-240000 = 480000"


Company B

"(70000+140000+165000)-150000=225000"


(2) current ratio= "\frac{Current \space assets}{Current \space liabilities}"

Company A

"\frac{720000}{240000}=3"


Company B

"\frac{375000}{150000}=2.5"


(3) Quick ratio= "\frac{Quick \space assets }{Quick \space liabilities}"

Company A

"\frac{36000+180000}{240000}=0.9"


Company B

"\frac{70000+140000}{150000}=1.4"


(4) Number of times inventory turned over during the year and the average number of days required to turn over the inventory

= "\frac{Cost \space of \space goods \space sold }{Inventory}"

Company A

"\frac{1260000}{504000}=2.5"


Number of days

"\frac{365}{2.5}=146"


Company B

"\frac{825000}{165000}=5"


Number of days

"\frac{365}{5}=73"


(5) Number of times accounts receivables turned over during the year and the average numbers of days required to collect accounts receivables

= "\frac{Net \space credit \space sales}{Net \space a/c \space receivable}"

Company A

"\frac{1440000}{180000}=8"


Number of days

"\frac{365}{8}=46"


Company B

"\frac{1190000}{140000}=8.5"


Number of days

"\frac{365}{8.5}=43"


(6) operating cycle = Inventory period + a/cs receivable period.

Company A

"146+46=192"

Company B

"73+43=116"